User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 20th July

Chair: Bryn

Present: Mas, Lucas, Ross, Patrick, Tiege, Jamie, Jesse

Apologies: Liz, Arwin, Lucia, Jonny


General Issues:

  • Arwin can't make every second meeting this semester - is there a better time for everyone?
    • Doodle poll to find a new time.
  • Any fix for the all-sky cam at Katherine?
    • Not sure if software or hardware issue - will be looked at by Jamie.
  • 26 meter being used for some geometry and austral experiments.
  • hb getting new dbbc - potentially more fiddly, will be operational in a few months' time.

Handover Issues:

  • Issues with Ho for aua023 - mark5 issues caused by bad prc file?
    • Schedules have an incorrect setting for the mk4 rack/mk5. Can improve files from the scheduler - Moving away from using the mk4 rack.
    • r1802 will be running on both mk5 and mk4 rack at same time to test moving over from the mk4 rack. No more angry red lights. If this works, we will be decommissioning the mk4 rack.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_07_20.txt · Last modified: 2017/07/20 03:20 by Jonathan Rogers